OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. B(6, 2), m = -1/2

OpenStudy (lgbasallote):

use point slope form. you know how to do that?

OpenStudy (anonymous):

um no ina serious rush juss need the answer thnx

OpenStudy (lgbasallote):

ok... 2y-4 = -x - 6 x+2y - 2 =0

OpenStudy (anonymous):

thanks